OpenStudy (anonymous):

Shayla is baking muffins that require 1/2 cup vegetable oil. She only has 1/3 cup. How much more does she need? (Please explain)

OpenStudy (xishem):

Shayla has some amount (1/3 cup) and needs x more cups to have 1/3 cup:\[\frac{1}{3}+x=\frac{1}{2}\]

OpenStudy (anonymous):

what is \(\frac{1}{2}-\frac{1}{3}\) ?

OpenStudy (anonymous):

@satellite73 1/6 Thanks!

OpenStudy (lovelyanna):

1/2 - 1/3 = 1/6